# Channel Estimation for Beyond Diagonal RIS via Tensor Decomposition

### v1

@@ Abstract
This paper addresses the channel estimation problem for beyond diagonal reconfigurable intelligent surface (BD-RIS) from a tensor decomposition perspective. We first show that the received pilot signals can be arranged as a three-way tensor, allowing us to recast the cascaded channel estimation problem as a block Tucker decomposition problem that yields decoupled estimates for the involved channel matrices while offering a substantial performance gain over the conventional (matrix-based) least squares (LS) estimation method. More specifically, we develop two solutions to solve the problem. The first one is a closed-form solution that extracts the channel estimates via a block Tucker Kronecker factorization (BTKF), which boils down to solving a set of parallel rank-one matrix approximation problems. Exploiting such a low-rank property yields a noise rejection gain compared to the standard LS estimation scheme while allowing the two involved channels to be estimated separately. The second solution is based on a block Tucker alternating least squares (BTALS) algorithm that directly estimates the involved channel matrices using an iterative estimation procedure. We discuss the uniqueness and identifiability issues and their implications for training design. We also propose a tensor-based design of the BD-RIS training tensor for each algorithm that ensures unique decoupled channel estimates under trivial scaling ambiguities. Our numerical results shed light on the tradeoffs offered by BTKF and BTALS methods. Specifically, while the first enjoys fast and parallel extraction of the channel estimates in closed form, the second has a more flexible training design, allowing for a significantly reduced training overhead compared to the state-of-the-art LS method.
@@ I Introduction
As a new advance of conventional reconfigurable intelligent surface (RIS) techniques with diagonal phase shift matrices [1, 2], beyond diagonal (BD) RIS has been recently proposed and theoretically proved to achieve enhanced channel gain and enlarged coverage [3, 4, 5]. The benefits of BD-RIS are enabled by interconnecting elements via additional tunable components to mathematically generate scattering matrices with nonzero off-diagonal entries, increasing flexibility to manipulate waves. The fundamental modeling and architecture (group/fully-connected) design of BD-RIS based on the circuit topology has been first studied in [6]. Following [6], other architectures (forest/tree-connected) with reduced circuit complexity yet satisfactory performance have been proposed based on the graph theory [7]. In [8], a frequency-dependent model is proposed to optimize a BD-RIS architecture for multi-band MIMO networks. Meanwhile, to enlarge the coverage, BD-RIS with hybrid transmissive and reflective mode and multi-sector mode have been proposed based on the antenna array arrangement [3].
@@
It is worth noting that the enhanced performance of BD-RIS architectures and modes depends highly on the accuracy of the channel state information (CSI). However, it is difficult to effectively and efficiently acquire the CSI for BD-RIS-aided wireless systems for the following reasons. First, since the BD-RIS is nearly passive without the ability to sense signals, a straightforward strategy is to estimate the combined channel constructed by the transmitter-RIS and RIS-user channels, as well studied in conventional RIS literature [9, 10, 11], [12, 13]. This strategy relies on the pre-design of BD-RIS for pilot training, while each BD-RIS architecture leads to unique mathematical constraints of the scattering matrix, which indicates that the design for conventional RISs does not work for BD-RIS architectures. Second, each BD-RIS architecture leads to unique constructions of the combined channel with increasing dimensions, which requires additional training overhead to obtain the CSI. To solve the above two challenges, one recent work [5] has proposed a closed-form solution based on the least squares (LS) estimation to pre-design the BD-RIS with group/fully-connected architectures. Nevertheless, there are two limitations in [5]. First, the combined CSI is obtained at the cost of a large training overhead, which grows heavily with the circuit complexity of BD-RIS architectures. Second, the built-in block Kronecker structure of the combined channel is ignored, which could be further exploited to facilitate channel estimation. Thus, estimating the BD-RIS-aided channels with high accuracy and low training overhead remains an important yet challenging open problem.
@@
To address the above limitations, this paper studies channel estimation for BD-RIS via a tensor decomposition perspective. Tensor methods have been successfully exploited in different knowledge areas, including signal processing and machine learning, as well as in wireless communications and multi-sensor processing (see [14, 15, 16, 17, 18] and references therein). Recently, tensor modeling has also found applications in RIS-assisted communications [19, 20, 21, 22, 23]. Here, we derive efficient channel estimation methods for BD-RIS that leverage the tensor decomposition structure of the received pilots to provide accurate CSI acquisition of each individual channel while operating at low training overheads. Specifically, we first show that the combined channel for BD-RIS can be formulated as a (block) Tucker tensor decomposition. Then, decoupled estimates for the involved channel matrices can be obtained by exploiting the different unfoldings of the received pilot tensor. We develop two solutions to solve this problem. The first one is a closed-form solution that extracts the individual channel estimates via a block Tucker Kronecker factorization (BTKF), which boils down to solving a set of parallel rank-one matrix approximation problems. The second one is based on a block Tucker alternating least squares (BTALS) algorithm that directly estimates the involved channels using an iterative estimation procedure. The proposed algorithms offer substantial performance gains over conventional LS estimation while operating at much lower training overheads by capitalizing on the tensor signal structure.
@@
First, we link the channel estimation problem for BD-RIS to a tensor decomposition problem. Specifically, we show that the received pilot signals can be organized as a three-dimensional (3D) array or a third-order tensor that follows a (block) Tucker decomposition model. In addition, we discuss the implications of the specific BD-RIS architecture to the resulting tensor decomposition structure.
@@
Third, we discuss the trade-offs involving the channel estimation methods and their implications for the training design. On the one hand, we show that the BTKF method effectively exploits the inherent Kronecker structure of the combined channel to provide a more accurate reconstruction than the reference LS method, thanks to the noise rejection property of the channel separation step. On the other hand, by efficiently exploiting the tensor decomposition structure of the received pilots, BTALS yields decoupled channel estimates with a much lower training overhead, which can be orders of magnitude smaller than that of LS and BTKF methods. Such savings in training resources offered by BTALS are even more pronounced for BD-RIS configurations with higher levels of couplings among the scattering elements.
@@
Finally, simulation results show the superiority of the proposed BTKF and BTALS algorithms compared to the baseline LS estimator while highlighting the trade-offs involving these methods in terms of normalized mean square error (NMSE) performance, required training overhead, and computational complexity. In particular, we show that BTKF and BTALS offer performance gains over the reference LS method thanks to decoupling the estimated channel matrices, yielding a more accurate reconstruction of the combined channel. For a group-connected BD-RIS architecture with $Q$ groups, the BTKF method achieves channel separation by solving a set of $Q$ rank-one matrix approximations in parallel. On the other hand, the BTALS method directly estimates the involved channel matrices by intertwining the estimation of the transmitter-RIS and RIS-receiver channels in an iterative way using an alternating least squares procedure. Our results also show that the NMSE performance of the estimated channels with group-connected BD-RIS architectures is the same as that with conventional RIS in some scenarios.
@@ II Tensor Prerequisites
In this section, we provide the useful notations and main operators used in this papers as well an overview of the Tucker decomposition, in which will be used to develop the proposed channel estimation algorithm.
@@ II-A Notation and properties
We also use the following property of the Kroneker product
@@ II-B Slices and unfoldings
For convenience, we can also refer to the unfolding operation as $\big{[}\bm{\mathcal{Y}}\big{]}_{\left(n\right)}=\textrm{unfold}(\bm{\mathcal{Y}},n)$, $n=1,2,3$. The $n$-mode product, denoted as “×$\times_{n}$”, defines the multiplication between a tensor $\bm{\mathcal{Y}}$ and a matrix $\bm{A}$, leading to a tensor $\bm{\mathcal{Z}}$ with compatible dimensions, i.e., $\bm{\mathcal{Z}}=\bm{\mathcal{Y}}\times_{n}\bm{A}$. It can be computed by pre-multiplying the $n$-mode unfolding of $\bm{\mathcal{Y}}$ by the matrix $\bm{A}$, i.e., $[\bm{\mathcal{Z}}]_{(n)}=\bm{A}[\bm{\mathcal{Y}}]_{(n)}$. For example, the 1-mode product between $\bm{\mathcal{Y}}\in\mathbb{C}^{I\times J\times K}$ and $\bm{A}\in\mathbb{C}^{L\times I}$ yields $\bm{\mathcal{Z}}=\bm{\mathcal{Y}}\times_{1}\bm{A}\in\mathbb{C}^{L\times J\times K}$. It can be computed by $[\bm{\mathcal{Z}}]_{(1)}=\bm{A}[\bm{\mathcal{Y}}]_{(1)}\in\mathbb{C}^{L\times JK}$.
@@ III System Model
Let us consider a multiple-input multiple-output (MIMO) system assisted by a beyond diagonal reconfigurable intelligent surface (BD-RIS), as illustrated in Fig. 1, where the transmitter and the receiver are equipped with $M_{T}$ and $M_{R}$ antennas, respectively, and a BD-RIS with $N$ elements. For simplification, we assume that the direct link between the transmitter and the receiver is blocked.
@@ III-A Signal and channel models
Assuming that the transmitter sends orthogonal pilot sequences, which requires $T\geq M_{T}$, and after matched-filtering using the known pilots, we get
@@
where $\bm{Y}_{k}\doteq\bar{\bm{Y}}_{k}\bm{X}^{\text{H}}$ and $\bm{B}_{k}=\bar{\bm{B}}_{k}\bm{X}^{\text{H}}$ are the filtered pilot signals and noise. Considering a group-connected BD-RIS architecture [6], the $N$ reflecting elements are divided into $Q$ groups, each with $\bar{N}$ elements connected to each other, i.e., $N=\bar{N}\cdot Q$. In this case, the scattering matrix is expressed as $\bm{S}_{k}=\text{blkdiag}(\bm{S}^{(1)}_{k},\ldots\bm{S}^{(Q)}_{k})\in\mathbb{C}^{N\times N}$, where the $q$-th matrix $\bm{S}^{(q)}_{k}\in\mathbb{C}^{\bar{N}\times\bar{N}}$ satisfies $\bm{S}^{(q)\text{H}}_{k}\bm{S}^{(q)}_{k}=\bm{I}_{\bar{N}}$. Hence, (13) translates into a sum of $Q$ blocks
@@ III-B Least squares channel estimation
$\bm{H}\hskip 1.42271pt|\otimes|\hskip 1.42271pt\bm{G}=\big{[}\bm{H}^{(1)}\otimes\bm{G}^{(1)},\ldots,\bm{H}^{(Q)}\otimes\bm{G}^{(Q)}\big{]}\in\mathbb{C}^{M_{R}M_{T}\times N},$
is the combined block-Kronecker-structured MIMO channel matrix that concatenates the combined channels associated with the $Q$ BD-RIS groups. Defining $\bm{T}\doteq\bm{H}\hskip 1.42271pt|\otimes|\hskip 1.42271pt\bm{G}$ and $\bm{\bar{S}}=[\text{vec}\big{(}\bm{\bar{S}}_{1}\big{)},\ldots,\text{vec}\big{(}\bm{\bar{S}}_{K}\big{)}]^{\text{T}}\in\mathbb{C}^{K\times\bar{N}^{2}Q}$ and collecting the received signal over $K$ blocks we have
@@
where $\bm{B}$ is the corresponding noise term. An estimate of the combined channel $\bm{T}$ can be obtained by right-filtering using the known BD-RIS training matrix, i.e., $\hat{\bm{T}}=\bm{Y}(\bm{\bar{S}}^{\text{T}})^{\dagger}$ as a solution to the following least squares (LS) problem
@@
where $\hat{\bm{T}}\approx\hat{\bm{H}}\hskip 1.42271pt|\otimes|\hskip 1.42271pt\hat{\bm{G}}$ is an estimate of the combined channel. Note that assuming an orthogonal design for BD-RIS training matrix $\bm{\bar{S}}\in\mathbb{C}^{K\times\bar{N}^{2}Q}$, the solution is found by simplified matched filtering [5]. In this case, the estimate of the combined channel can also be found as $\hat{\bm{T}}=\bm{Y}\bm{\bar{S}}^{\ast}$. The LS solution requires $K\geq\bar{N}^{2}Q$ to ensure a unique estimation of the combined channel. This constraint may be too restrictive, especially for a moderate number of scattering elements, due to the quadratic dependency on the number of connected BD-RIS elements in each group.
III-C BD-RIS matrix design and motivation
Although the design proposed in [5] is optimal in the mean square error (MSE) sense, LS channel estimation based on (18) ignores the built-in block Kronecker structure of the effective MIMO channel since only an estimate of the “combined” channel is obtained. Hence, it requires a large training overhead, which has a quadratic growth with the group size $\bar{N}$. As discussed in the next sections, the block Kronecker product structure linking the involved channel matrices can be efficiently exploited, allowing us to obtain enhanced channel estimates compared to the baseline LS solution. Additionally, it turns out that the received signal in (14) can be recast using a tensor modeling approach. This is possible by reformulating the training design and rebuilding the BD-RIS training structure as a third-order tensor. Then, by resorting to tensor decomposition algorithms and capitalizing on their intrinsic uniqueness properties, decoupled estimates of the individual channels $\bm{G}$ and $\bm{H}$ can be obtained with a significantly reduced training overhead and improved accuracy. Additionally, as will be shown later (Section VI), the proposed BD-RIS training design is flexible to allow operation under more challenging system setups with $K<<\bar{N}^{2}Q$, which have not yet been considered in the literature.

### v2

@@ Abstract
This paper addresses the channel estimation problem for beyond diagonal reconfigurable intelligent surface (BD-RIS) from a tensor decomposition perspective. We first show that the received pilot signals can be arranged as a three-way tensor, allowing us to recast the cascaded channel estimation problem as a block Tucker decomposition problem that yields decoupled estimates for the involved channel matrices while offering a substantial performance gain over the conventional (matrix-based) least squares (LS) estimation method. More specifically, we develop two solutions to solve the problem. The first is a closed-form solution that extracts the channel estimates via a block Tucker Kronecker factorization (BTKF), which boils down to solving a set of parallel rank-one matrix approximation problems. Exploiting such a low-rank property yields a noise rejection gain compared to the standard LS estimation scheme, allowing the two involved channels to be estimated separately. The second solution is based on a block Tucker alternating least squares (BTALS) algorithm that directly estimates the involved channel matrices using an iterative estimation procedure. We discuss the uniqueness and identifiability issues and their implications for training design. We also propose a tensor-based design of the BD-RIS training tensor for each algorithm that ensures unique decoupled channel estimates under trivial scaling ambiguities. Our numerical results shed light on the tradeoffs offered by BTKF and BTALS methods. Specifically, while the first enjoys fast and parallel extraction of the channel estimates in closed form, the second has a more flexible training design, allowing for a significantly reduced training overhead compared to the state-of-the-art LS method.
@@ I Introduction
As a new advance of conventional reconfigurable intelligent surface (RIS) techniques with diagonal phase shift matrices [1, 2], beyond diagonal (BD) RIS has been recently proposed and theoretically proved to achieve enhanced channel gain and enlarged coverage [3, 4, 5]. The benefits of BD-RIS are enabled by interconnecting elements via additional tunable components to mathematically generate scattering matrices with nonzero off-diagonal entries, increasing flexibility to manipulate waves. The fundamental modeling and architecture (group/fully connected) design of BD-RIS based on the circuit topology has been first studied in [6]. Following [6], other architectures (forest/tree-connected) with reduced circuit complexity yet satisfactory performance have been proposed based on the graph theory [7]. In [8], a frequency-dependent model is proposed to optimize a BD-RIS architecture for multi-band MIMO networks. Meanwhile, BD-RIS with hybrid transmissive, reflective, and multi-sector modes have been proposed to enlarge the coverage, based on the antenna array arrangement [3].
@@
It is worth noting that the enhanced performance of BD-RIS architectures and modes depends highly on the accuracy of the channel state information (CSI). However, it is difficult to effectively and efficiently acquire the CSI for BD-RIS-aided wireless systems for the following reasons. First, since the BD-RIS is nearly passive without the ability to sense signals, a straightforward strategy is to estimate the composite channel constructed by the transmitter-RIS and RIS-user channels, as well studied in conventional RIS literature [9, 10, 11], [12, 13]. This strategy relies on the pre-design of BD-RIS for pilot training. At the same time, each BD-RIS architecture leads to unique mathematical constraints of the scattering matrix, which indicates that the design for conventional RISs does not work for BD-RIS architectures. Second, each BD-RIS architecture leads to unique constructions of the composite channel with increasing dimensions, which requires additional training overhead to obtain the CSI. To solve the above two challenges, one recent work [5] has proposed a closed-form solution based on the least squares (LS) estimation to pre-design the BD-RIS with group/fully-connected architectures. Nevertheless, there are two limitations in [5]. First, the composite CSI is obtained at the cost of a large training overhead, which grows heavily with the circuit complexity of BD-RIS architectures. Second, the built-in block Kronecker structure of the composite channel is ignored, which could be further exploited to facilitate channel estimation. Thus, estimating the BD-RIS-aided channels with high accuracy and low training overhead remains an essential yet challenging open problem.
To address the above limitations, this paper studies channel estimation for BD-RIS via a tensor decomposition perspective.
Tensor modeling has been widely adopted in signal processing for wireless communications due to its ability to manipulate and efficiently exploit the intrinsic multidimensional structure of wireless communication signals and channels (usual dimensions are space, time, frequency, polarization, coding, etc.) [14, 15, 16, 17, 18, 19, 20, 21, 22]. In the case of reconfigurable intelligent surface (RIS), tensor decompositions have proven to provide effective solutions to channel estimation and beamforming design [23, 24, 25, 26, 27]. This is due to the ability of tensors to decouple individual estimates of the involved channels using decompositions such as the PARAllel FACtors (PARAFAC) and Tucker decompositions under more flexible conditions than competing methods. Tensor-based algorithms have recently been used in integrated sensing and communication (ISAC) systems [28, 29, 30]. The work of [28] models the received sensing signal using a PARAFAC decomposition, and the authors propose an alternating least squares (ALS) algorithm to estimate the sensing factors. On the other hand, in [29] the authors exploit the sparsity of the THz channels and model the signal using a PARAFAC tensor while proposing an ALS based algorithm to estimate the channel parameters. In [30], a PARAFAC tensor model is proposed for a channel estimation framework in MIMO-OFDM with dynamic metasurface antennas (DMA). Different from [28, 29, 30], the proposed approach focuses on channel estimation for beyond diagonal reconfigurable intelligent surface (BD-RIS) assisted multiple-input multiple-output (MIMO) systems, by employing a tensor-based model which allows for a decoupled estimate of the involved channels.
It is worth noting that in the case of single-connected (diagonal) RIS, the solution proposed in [23] showed that the individual channel estimation problem could be solved using the PARAFAC decomposition, allowing us to find decoupled channel estimates of the involved channels, while providing a refined estimate of the composite TX-RIS-RX channel. However, individual channel estimation methods valid for traditional single-connected RIS do not apply to BD-RIS due to its more general (and complex) inter-element connection structure and design constraints. More specifically, the PARAFAC-based individual channel estimation method of [23] cannot capture the algebraic tensor structure of the received pilots, which requires rethinking channel training in two aspects. First, an adequate tensor decomposition approach is necessary to obtain decoupled channel estimates from the received pilot signals, and second, BD-RIS training structures are equivalent to third-order tensors (as opposed to diagonal RIS training structures modeled as matrices).
@@
This paper shows that the composite channel for BD-RIS can be formulated as a (block) Tucker tensor decomposition. Decoupled estimates for the involved channel matrices can be obtained by exploiting the different unfoldings of the received pilot tensor. We develop two solutions to solve this problem. The first is a closed-form solution that extracts the individual channel estimates via a block Tucker Kronecker factorization (BTKF), which boils down to solving a set of parallel rank-one matrix approximation problems. The second one is based on a block Tucker alternating least squares (BTALS) algorithm that directly estimates the involved channels using an iterative estimation procedure. The proposed algorithms offer substantial performance gains over conventional LS estimation while operating at much lower training overheads by capitalizing on the tensor signal structure.
@@
First, we link the channel estimation problem for BD-RIS to a tensor decomposition problem. Specifically, we show that the received pilot signals can be organized as a three-dimensional (3D) array or a third-order tensor that follows a (block) Tucker decomposition model. In addition, we discuss the implications of the specific BD-RIS architecture on the resulting tensor decomposition structure.
@@
Third, we discuss the trade-offs involving the channel estimation methods and their implications for the training design. On the one hand, we show that the BTKF method effectively exploits the inherent Kronecker structure of the composite channel to provide a more accurate reconstruction than the reference LS method, thanks to the noise rejection property of the channel separation step. On the other hand, by efficiently exploiting the tensor decomposition structure of the received pilots, BTALS yields decoupled channel estimates with a much lower training overhead, which can be orders of magnitude smaller than that of LS and BTKF methods. Such savings in training resources offered by BTALS are even more pronounced for BD-RIS configurations with higher levels of couplings among the scattering elements.
@@
Finally, simulation results show the superiority of the proposed BTKF and BTALS algorithms compared to the baseline LS estimator while highlighting the trade-offs involving these methods in terms of normalized mean square error (NMSE) performance, required training overhead, and computational complexity. In particular, we show that BTKF and BTALS offer performance gains over the reference LS method thanks to decoupling the estimated channel matrices, yielding a more accurate reconstruction of the composite channel. For a group-connected BD-RIS architecture with $Q$ groups, the BTKF method achieves channel separation by solving a set of $Q$ rank-one matrix approximations in parallel. On the other hand, the BTALS method directly estimates the involved channel matrices by intertwining the estimation of the transmitter-RIS and RIS-receiver channels in an iterative way using an alternating least squares procedure. Our results also show that the NMSE performance of the estimated channels with group-connected BD-RIS architectures is the same as that with conventional RIS in some scenarios.
@@ II Tensor Prerequisites
In this section, we provide the proper notations and main operators used in this paper and an overview of the Tucker decomposition, which will be used to develop the proposed channel estimation algorithm.
@@ II-A Notation and properties
We also use the following property of the Kronecker product
@@ II-B Slices and unfoldings
For convenience, we can also refer to the unfolding operation as $\big{[}\bm{\mathcal{Y}}\big{]}_{\left(n\right)}=\textrm{unfold}(\bm{\mathcal{Y}},n)$, $n=1,2,3$. The $n$-mode product, denoted as “×$\times_{n}$”, defines the multiplication between a tensor $\bm{\mathcal{Y}}$ and a matrix $\bm{A}$, leading to a tensor $\bm{\mathcal{Z}}$ with compatible dimensions, i.e., $\bm{\mathcal{Z}}=\bm{\mathcal{Y}}\times_{n}\bm{A}$. It can be computed by pre-multiplying the $n$-mode unfolding of $\bm{\mathcal{Y}}$ by the matrix $\bm{A}$, i.e., $[\bm{\mathcal{Z}}]_{(n)}=\bm{A}[\bm{\mathcal{Y}}]_{(n)}$. For example, the 1-mode product between $\bm{\mathcal{Y}}\in\mathbb{C}^{I\times J\times K}$ and $\bm{A}\in\mathbb{C}^{L\times I}$ yields $\bm{\mathcal{Z}}=\bm{\mathcal{Y}}\times_{1}\bm{A}\in\mathbb{C}^{L\times J\times K}$. It can be computed by $[\bm{\mathcal{Z}}]_{(1)}=\bm{A}[\bm{\mathcal{Y}}]_{(1)}\in\mathbb{C}^{L\times JK}$. We refer the interested reader to [20] for an overview.
@@ III System Model
Let us consider a MIMO system assisted by a BD-RIS, as illustrated in Fig. 1, where the transmitter and the receiver are equipped with $M_{T}$ and $M_{R}$ antennas, respectively, and a BD-RIS with $N$ elements. For simplification, we assume that the direct link between the transmitter and the receiver is blocked.
@@ III-A Signal and channel models
Assuming that the transmitter sends orthogonal pilot sequences, which requires $T\geq M_{T}$, and after matched-filtering using the known pilots, we get [^2] According to [34], when taking into account the mutual coupling between BD-RIS elements, the term $\bm{S}_{k}$ in (13) is replaced by $(\bm{S}_{k}^{-1}-\bm{S}_{II})^{-1}$, where $\bm{S}_{II}$ denotes the mutual coupling matrix. This means the whole channel estimation scheme will still work by regarding the term $(\bm{S}_{k}^{-1}-\bm{S}_{II})^{-1}$ as a whole, while the design of $\bm{S}_{k}$ should be carefully re-considered to capture the impact induced by $\bm{S}_{II}$. In this work, we have neglected the coupling effects in BD-RIS since the study of this aspect is out of the scope of this paper. This assumption can be achieved when 1) the inter-element spacing is sufficiently large (at least half-wavelength), and 2) the antenna elements in BD-RIS are perfectly matched [35]. This allows us to have $\bm{S}_{II}=\bm{0}$, leading to a linear channel model widely adopted in the RIS literature.
@@
where $\bm{Y}_{k}=\bar{\bm{Y}}_{k}\bm{X}^{\text{H}}$ and $\bm{B}_{k}=\bar{\bm{B}}_{k}\bm{X}^{\text{H}}$ are the filtered pilot signals and noise. Considering a group-connected BD-RIS architecture [6], the $N$ reflecting elements are divided into $Q$ groups, each with $\bar{N}$ elements connected, i.e., $N=\bar{N}\cdot Q$. In this case, the scattering matrix is expressed as $\bm{S}_{k}=\text{blkdiag}(\bm{S}^{(1)}_{k},\ldots\bm{S}^{(Q)}_{k})\in\mathbb{C}^{N\times N}$, where the $q$-th matrix $\bm{S}^{(q)}_{k}\in\mathbb{C}^{\bar{N}\times\bar{N}}$ satisfies $\bm{S}^{(q)\text{H}}_{k}\bm{S}^{(q)}_{k}=\bm{I}_{\bar{N}}$. Hence, (13) translates into a sum of $Q$ blocks
@@ III-B Least squares channel estimation
$\bm{H}\hskip 1.42271pt|\otimes|\hskip 1.42271pt\bm{G}=\big{[}\bm{H}^{(1)}\otimes\bm{G}^{(1)},\ldots,\bm{H}^{(Q)}\otimes\bm{G}^{(Q)}\big{]}\in\mathbb{C}^{M_{R}M_{T}\times\bar{N}^{2}Q},$
is the composite block-Kronecker-structured MIMO channel matrix that concatenates the composite channels associated with the $Q$ BD-RIS groups. Defining $\bm{T}\doteq\bm{H}\hskip 1.42271pt|\otimes|\hskip 1.42271pt\bm{G}$ and $\bm{\bar{S}}=[\text{vec}\big{(}\bm{\bar{S}}_{1}\big{)},\ldots,\text{vec}\big{(}\bm{\bar{S}}_{K}\big{)}]^{\text{T}}\in\mathbb{C}^{K\times\bar{N}^{2}Q}$ and collecting the received signal over $K$ blocks we have
@@
where $\bm{B}$ is the corresponding noise term. An estimate of the composite channel $\bm{T}$ can be obtained by right-filtering using the known BD-RIS training matrix, i.e., $\hat{\bm{T}}=\bm{Y}(\bm{\bar{S}}^{\text{T}})^{\dagger}$ as a solution to the following least squares (LS) problem
@@
where $\hat{\bm{T}}\approx\hat{\bm{H}}\hskip 1.42271pt|\otimes|\hskip 1.42271pt\hat{\bm{G}}$ is an estimate of the composite channel. Note that assuming an orthogonal design for BD-RIS training matrix $\bm{\bar{S}}\in\mathbb{C}^{K\times\bar{N}^{2}Q}$, the solution is found by simplified matched filtering [5]. In this case, the estimate of the composite channel can also be found as $\hat{\bm{T}}=\bm{Y}\bm{\bar{S}}^{\ast}$. The LS solution requires $K\geq\bar{N}^{2}Q$ to ensure a unique estimation of the composite channel. This constraint may be too restrictive, especially for a moderate number of scattering elements, due to the quadratic dependency on the number of connected BD-RIS elements in each group.
III-C Motivation
Although the design proposed in [5] is optimal in the mean square error (MSE) sense, LS channel estimation based on (18) ignores the built-in block Kronecker structure of the effective MIMO channel since only an estimate of the “composite” channel is obtained. Hence, it requires a significant training overhead with a quadratic growth with the group size $\bar{N}$. As discussed in the following sections, the block Kronecker product structure linking the involved channel matrices can be efficiently exploited, allowing us to obtain enhanced channel estimates compared to the baseline LS solution. Additionally, it turns out that the received signal in (14) can be recast using a tensor modeling approach. This is possible by reformulating the training design and rebuilding the BD-RIS training structure as a third-order tensor. Then, by resorting to tensor decomposition algorithms and capitalizing on their intrinsic uniqueness properties, decoupled estimates of the individual channels $\bm{G}$ and $\bm{H}$ can be obtained with a significantly reduced training overhead and improved accuracy. Additionally, as will be shown later (Section VI), the proposed BD-RIS training design is flexible to allow operation under more challenging system setups with $K<<\bar{N}^{2}Q$, which have not yet been considered in the literature.